OpenStudy (waheguru):

How to solve this? 70 = -5t^2+20t+50? the answer is 2 but i am getting 1 ?

OpenStudy (***[isuru]***):

hi, divide the whole by 5 first... 14 = -t^2 + 4t + 10 now... t^2 -4t + 4 = 0 now this is a complete square.. (t - 2)^2 = 0 now did u get it ?

OpenStudy (waheguru):

I get it but I am trying to find the x value for teh vertex when the y value is 70

OpenStudy (waheguru):

So How can I solve for t algebraically?

OpenStudy (***[isuru]***):

show ur steps.... then we can tell ya wt's wrong with them... that way u'll understand it better

OpenStudy (waheguru):

I dont even know how to start?

OpenStudy (***[isuru]***):

but ya said that ya get 1 instead of 2 ?

OpenStudy (waheguru):



OpenStudy (anonymous):

70=-5(t^2-4t-10)

OpenStudy (***[isuru]***):

ya basically wants to find the value of t when y = 70 am I right ?

OpenStudy (waheguru):

yea

OpenStudy (waheguru):

How can I do that?

OpenStudy (waheguru):

I got it, I have to complete the square which gives me -5(t-2)^2-20 and then I can set t-2 = 0 and then t = 2 which is the correct answer

OpenStudy (***[isuru]***):

yep!! ya got it man

OpenStudy (anonymous):

i was just about to suggest that!